### v1

@@ Appendix 0.E Open-Ended Object Detection
In this experiment, we compare the effectiveness of a describe-then-localize method (specifically LLAVA 1.5 [32] 7B + OWL-ViTv2 [34]) with a localize-then-describe method (specially FlexCap + OWL-ViTv2) for the task of open-ended object detection. For the describe-then-localize method, we generate a list of objects using the following prompt: "Describe object names and regions in this image." We extract the nouns and use them as text queries for OWL-ViTv2. For the localize-then-describe approach, we take the top 128 objects ranked by objectness score by OWL-ViTv2 and describe them with different length prompts. In Fig. 5 in the main paper, we compare recall of the ground-truth regions annotated in the Visual Genome dataset for both these approaches. We find that while describe-then-localize with LLAVA can be an effective approach for finding large objects, localize-then-describe with FlexCap is significantly better at medium and small sized objects and marginally better for large-sized objects.
Appendix 0.F Visual Dialog
@@
FlexCap-LLM can be used for the task of visual dialog [10]. We first caption all the objects in the image using FlexCap and OWL-ViTv2. We retain the top 128 boxes according to the objectness score from OWL-ViTv2 and describe each box using FlexCap. See Fig. 1 and Fig. 4 in the main paper for more details. Once the image has been represented as the list of objects in text, we can interact with a LLM by providing the conversation turns as additional context for each query to the LLM. We show some examples of conversations with the FlexCap-LLM system in Figure 15. Note how the model is able to read text in the image in the leftmost figure, recognize material in the middle figure, and localize objects of interest in the rightmost figure. As we compute the object captions only once at the beginning of the conversation, there is no additional overhead of querying a large VLM for each additional turn in the conversation. We show more examples of this on the webpage in the FlexCap-LLM section of https://flex-cap.github.io.
@@
Appendix 0.G LLM Prompts
We use the following prompts for the LLM in the question-answering experiments.
OK-VQA and GQA.
@@ Appendix 0.G LLM Prompts
  objects_description += f"[{captions}] [{cx}, {cy}, {w}, {h}],"
@@
for captions, (cx, cy, w, h), score in zip(object_captions, object_boxes, object_scores):
  objects_description += f"[{captions}] [{cx}, {cy}, {w}, {h}] [{score :.1f}],"
@@
    objects_description += f"[{captions}],"

### v2

@@ Appendix C Open-Ended Object Detection
In this experiment, we compare the effectiveness of a describe-then-localize method (specifically LLAVA 1.5 [33] 7B + OWL-ViTv2 [35]) with a localize-then-describe method (specially FlexCap + OWL-ViTv2) for the task of open-ended object detection. For the describe-then-localize method, we generate a list of objects using the following prompt:Describe object names and regions in this image. We extract the nouns and use them as text queries for OWL-ViTv2. For the localize-then-describe approach, we take the top 128 objects ranked by objectness score by OWL-ViTv2 and describe them with different length prompts. In Fig. 5 in the main paper, we compare recall of the ground-truth regions annotated in the Visual Genome dataset for both these approaches. We find that while describe-then-localize with LLAVA can be an effective approach for finding large objects, localize-then-describe with FlexCap is significantly better at medium and small sized objects and marginally better for large-sized objects. We show some typical examples of missed detections in Fig. 11.
Figure 12: Diverse captioning with Prefixes. FlexCap can be used to perform conditional captioning of images. We show three prefixes: "a", "the photo was taken in ", "notice" resulting in diverse captions for the same image. Note the input red bounding box is around the full image.
Appendix D Different prefixes with same image
In Fig. 12, we highlight how different prefixes can be used to generate diverse captions. Note how in column 1 the model correctly identifies the country even, possibly using the logo on the train. We find the prefix "Notice" leads to captions highlighting noteworthy aspects in an image something which the un-prefixed caption does not do.
Appendix E Visual Dialog
@@
FlexCap-LLM can be used for the task of visual dialog [12]. We first caption all the objects in the image using FlexCap and OWL-ViTv2. We retain the top 128 boxes according to the objectness score from OWL-ViTv2 and describe each box using FlexCap. See Fig. 1 and Fig. 4 in the main paper for more details. Once the image has been represented as the list of objects in text, we can interact with a LLM by providing the conversation turns as additional context for each query to the LLM. We show some examples of conversations with the FlexCap-LLM system in Figure 13. Note how the model is able to read text in the image in the leftmost figure, recognize material in the middle figure, and localize objects of interest in the rightmost figure. As we compute the object captions only once at the beginning of the conversation, there is no additional overhead of querying a large VLM for each additional turn in the conversation. We show more examples of this on the webpage in the FlexCap-LLM section of index.html.
@@
Appendix F Effect of Length on FlexCapLLM
In this experiment, we study the effect of generating captions of different lengths on downstream visual question answering tasks. We evaluate this by using captions of increasing lengths on the val split of the VQAv2 dataset. We report the results of this experiment in Table 6. We find that increasing the length of generated captions results in better VQAv2 and GQA performance.
Maximum Length of Captions
Dataset 1 2 4 8
VQAv2 (val split) 57.4 59.9 66.4 67.8
GQA (testdev-balanced split) 43.6 45.3 47.7 48.8
Table 6: Effect of caption lengths on FlexCap-LLM
Appendix G FlexCapLLM Details
In Fig. 14 we show the whole pipeline how FlexCap interfaces with LLMs. We convert an image into a list of localized texts. These texts are passed onto the LLM with a preamble indicating the task and the VQA task’s question. We use the following prompts for the LLM in the question-answering experiments.
VQAv2, OK-VQA and GQA.
@@ Appendix G FlexCapLLM Details
objects_description += f"{captions} [{cx}, {cy}, {w}, {h}],"
@@
for captions, (cx, cy, w, h) in zip(object_captions, object_boxes):
objects_description += f"{captions} [{cx}, {cy}, {w}, {h}],"
@@
objects_description += f"{captions},"